Problem - A sample of gallium arsenide (GaAs) is doped with 1010 silicon atoms per cm3


.


Ninety-five percent of the silicon atoms replace arsenic atoms and the remaining five percent


replace gallium atoms. T = 300 K. The intrinsic carrier concentration, ni


, is equal to 9 × 106 cm3


.


Calculate the electron and hole concentration as well as the position of the Fermi level.

1
Expert's answer
2022-02-03T09:03:39-0500

1)

"n_1=2\cdot 10^{16}," "p_1=5\cdot 10^3,"

"n_2=10^{16}," "p_2=10^4,"

2)

"n_1=8\cdot10^{15}," "p_1=1.25\cdot 10^4,"

"n_2=10^{16}," "p_2=10^4."